OpenStudy (anonymous):

Select the point that lies within the solution to the graph of y > –3x + 4. (–1, 7) (0, 3) (1, 3) (3, –6)

OpenStudy (callisto):

Put (x,y) into the equation and see if the inequality is true. For example, put (-1, 7) into y > –3x + 4. Left side = 7 Right side = -3(-1) + 4 = 3+4 = 7 = left side So, (-1, 7) is not a pair of solution. Can you do it now?